OpenStudy (anonymous):

Find the volume of the cylinder in terms of Pi. r=1.9 m h=11 m

OpenStudy (anonymous):

@ybarrap

OpenStudy (anonymous):

pi r squre h

OpenStudy (ybarrap):

 You need to find the volume of the figure above. Take the area of the bottom surface and multiply it by the height of the cylinder.

OpenStudy (anonymous):

3.14 * 1.9*1.9 *11

OpenStudy (ybarrap):



OpenStudy (anonymous):

11.3411494794591536 is what i got for area @ybarrap

OpenStudy (ybarrap):

You need to provide answer in terms of pi.

OpenStudy (ybarrap):

do you know what that means?

OpenStudy (mathmate):

@LetsLearn2000 "in terms of pi" means do not evaluate pi. Pi is a trancendental number meaning that a decimal representation is not exact. For example, a circle with a radius of 4 has an area of \(\pi 4^2\)=\(16\pi\), in terms of pi. so with r=1.9, h=11m, the volume in terms of pi is therefore V=\(\pi 1.9^2\times 11\)= \(39.71\pi \ m^3\)